# Cheating Stereo Matching in Full-scale: Physical Adversarial Attack against Binocular Depth Estimation in Autonomous Driving

### v2

@@ Introduction
Figure 1: Comparison between existing PAEs and our 3D texture-enabled merging attack PAEs in CARLA simulator (dosovitskiy2017carla). (a) Benign right and left camera images with the corresponding depth map; (b) Random texture; (c), (d) 2D adversarial patches from (liu2024physical) and (cheng2021towards), respectively, applied to 3D vehicle. (e) MDE Adversarial texture from (zheng2024physical); (f) Our 3D merging attack adversarial texture.
@@
Recent studies have revealed the vulnerabilities of deep neural perception models for Autonomous Driving (AD) against physical adversarial attacks (lovisotto2021slap; zhu2021can; cao2021invisible). These attacks rely on physically deployable artifacts, also known as Physical Adversarial Examples (PAEs) (eykholt2018robust; thys2019fooling), to affect the vision features captured by the neural networks and then manipulate their predictions. Currently, most PAEs are in the form of 2D patches that affect only a limited fraction of an object, and they often target only Monocular Depth Estimation (MDE) models (cheng2022physical; guesmi2023aparate; wong2021stereopagnosia; cheng2021towards). Consequently, these threats to AD might be readily bypassed in real-world 3D scenarios with the support of stereo-based Binocular Depth Estimation (BDE) (xu2020aanet; wang2019pseudo; li2019stereo).
@@
In fact, attacks on BDE driven by Stereo Matching (SM), a depth estimation method leveraging visual disparity, remain very limited. Sharply different from the PAEs for MDE, attacking SM models presents unique challenges due to its reliance on the physical 3D disparity between two images, rather than only on the 2D image pixels captured by a single camera; it is this uniqueness that invalidates most of MDE attacks under SM-BDE. Although a few proposals (liu2024physical; cheng2021towards) have explored attacks on stereo matching, they mainly rely on 2D patches overlaid in image space, which not only limits their effectiveness to narrow viewpoints, but also breaks the physical disparity consistency required by real-world systems, rendering them ineffective under physical deployment. Meanwhile, SM-BDE is increasingly adopted in AD, with dedicated datasets available to provide stereo data for training SM-BDE models (geiger2012we; mayer2016large). Therefore, it makes sense to consider full-scale 3D PAEs that are effective against SM-BDE, because only such PAEs can substantiate their realistic threat to real-world AD. To generate such PAE, one would need a global 3D camouflage texture to cover the whole surface of a target object, because such a texture can significantly increase the affected range within camera views and thus achieve a full-scale adversarial influence on the target.
@@
Adapting PAEs to SM-BDE may also enable more meaningful attacks. Existing attacks on depth estimation primarily focused on hiding objects by pushing their estimated depth to infinity (guesmi2023aparate; zheng2024physical; cheng2022physical). As a result, such attacks fail to really hide any objects, as the contour of the object remains, totally exposing the attackers’ intention. A more meaningful attack, as one may intuitively expect, should involves adjustable factors for depth fine-tuning, so that attackers can completely adapt the depth of different parts of the target object (including contours) into the surroundings, which we refer to as merging attack. Nonetheless, realizing the 3D PAE faces a great challenge under SM-BDE, as each pair of rendered images must precisely align with the physical geometry of stereo cameras in the real world. In addition to making PAEs viable under SM-BDE, a qualified design should also remain robust to varying viewpoints (e.g., position and heading) of the SM camera pair, as well as to diversified external environments (e.g., lighting and weather conditions).
@@
Additionally, our PAE optimization accounts for diverse environmental variations, including changes in distances, rotations lighting, and weather conditions. Figure 1 presents our merging attack PAEs compared to prior works (liu2024physical; cheng2021towards; zheng2024physical). The results demonstrate that, under real-world 3D stereo matching constraints, only our method remains effective. Our main contributions are summarized as follows:
@@
SM-BDE has received significant attention for its reliability and efficiency in depth estimation for AD (apollo; waymo; mobileye). Early works (zagoruyko2015; zbontar2016) leverage deep networks for separate feature extraction and matching. Subsequent advances include PSMNet (chang2018), which incorporates pyramid pooling and 3D CNNs; GA-Net (zhang2019ga), which introduces attention-based cost aggregation; and RAFT-Stereo (lipson2021raft), which uses recurrent units for multi-level refinement. Recent works explore cascaded recurrent networks, such as CREStereo (li2022practical), and focus on robust generalization and arbitrary-scale disparity, as demonstrated by AnyStereo (liang2024any). Given the rapid development of SM-BDE in AD, addressing its security has become critical.
@@
Adversarial attacks exploit vulnerabilities in DNN by introducing crafted perturbations to input data, causing the model to make incorrect predictions (goodfellow2014explaining; madry2017towards). PAEs have shown real-world effectiveness (kurakin2018adversarial; eykholt2018robust), sparking broad interest in safety-critical domains such as AD. They have been studied extensively across key AD tasks, including 3D and 2D object detection (tu2020physically; abdelfattah2021towards; cao2021invisible), object tracking (schmalfuss2022perturbation; ranjan2019attacking), object segmentation (zhu2021adversarial; xie2017adversarial), and MDE (cheng2022physical; guesmi2023aparate; wong2021stereopagnosia). Recent works (liu2024physical; cheng2021towards) also explored physical attacks on SM-BDE using 2D PAE but oversimplify stereo image relationships by assuming a direct shift between left and right views, neglecting stereo viewpoint differences. While this assumption may be efficient for digital 2D patch attacks, it fails to generalize to the complexities of real-world 3D scenarios, as shown in Figure 1. However, our 3D camouflage texture as PAE can conform to the parametric constraints of stereo cameras and achieve state-of-the-art attack effectiveness across diverse viewpoints on SM models.
@@
To ensure that the adversarial texture is both physically printable and visually smooth, we further incorporate two regularization terms: a total variation (TV) loss ($\mathcal{L}_{\text{tv}}$(mahendran2015understanding) that suppresses high-frequency noise, and a non-printability score (NPS) loss ($\mathcal{L}_{\text{nps}}$(sharif2016accessorize) that encourages the use of printer-reproducible colors. The final loss function is:
@@
We train our adversarial textures using the KITTI2015 dataset (geiger2012we), a widely adopted benchmark for AD that provides stereo image pairs. The training samples cover diverse vehicle poses, scales, and backgrounds, promoting generalization across real-world scenes. To enhance the robustness of adversarial textures under diverse environmental conditions, we adopt Expectation over Transformation (EoT) (athalye2018synthesizing) during training. Specifically, we randomly perturb the position of a point light source within the range [3,3][-3,3] meters along each axis, and uniformly sample ambient light intensity from $[0.3,0.9]$. To simulate challenging weather such as rain and fog, we additionally inject random Gaussian noise into the rendered images during training. For boundary depth extraction, the kernel size $p_{k}$ is set to $40$. We optimize the adversarial texture using a differentiable renderer and train for 100 epochs with the Adam optimizer (kingma2014adam), with an initial learning rate of 0.01, scheduled via cosine decay to a minimum of 1e41e-4. The loss coefficients are set to $\alpha=5$ and $\beta=0.1$. All experiments are performed on a single NVIDIA RTX A5000 GPU.
@@
We compare our method with the two most relevant physical attacks designed against SM-BDE: PASM (liu2024physical) and Adv-DM (cheng2021towards). In addition, we include two baselines textures: benign (neutral textures) and random (random noise). We evaluate all textures on five representative SM-BDE models: PSMNet (chang2018), GA-Net (zhang2019ga), RAFT-Stereo (lipson2021raft), CREStereo (li2022practical), and AnyStereo (liang2024any), covering diverse cost volume constructions and model designs.
@@
We evaluate the trained textures in the CARLA simulator (dosovitskiy2017carla) using a world-aligned stereo setup. Adversarial vehicles are placed at distances ranging from 3 to 30 meters and orientations between $0^{\circ}$ and $360^{\circ}$, under varying lighting and weather conditions in the “Town 10” map. Table 1 summarizes the evaluation results of our merging attack in comparison with state-of-the-art physical attacks across five representative SM-BDE models. All attacks are tested under identical camera baselines and environmental conditions in the CARLA simulator. Our adversarial texture consistently achieves a higher $\mathcal{E}_{\text{cover}}$, indicating a broader impact on predicted depth. This can be attributed to the larger effective surface coverage and better generalization across object orientations. Moreover, our approach attains the lowest $\mathcal{E}_{\text{blend}}$ among all methods, suggesting superior blending into the surrounding environment and reduced detectability in depth predictions, both key objectives of hiding attacks. Importantly, when integrated into a full commercial AD pipeline – Apollo, our attack leads to the highest $\mathcal{E}_{\text{sys}}$, indicating a significant increase in missed detection and path planning failures. This highlights the safety risk posed by our physically realizable texture.
@@
We conduct an ablation study to evaluate the effectiveness of the Stereo-Aligned 3D Rendering (SAR) module and the merging attack mechanism. We compare adversarial textures generated under four configurations: no module (None), SAR only, merging only, and the full method with both components. In the merging only setting, we define the main loss $\mathcal{L}_{\text{main}}$ as $\mathcal{L}_{\text{merge}}$ and reuse the same rendering parameters $k$ from the left view for the right. For None and SAR only, we replace $\mathcal{L}_{\text{merge}}$ with a standard hiding loss (zheng2024physical). As shown in Table 4, removing the physical constraint $\mathcal{L}_{\text{nps}}$ slightly improves performance due to the relaxation of printability limits. However, without enforcing stereo consistency, the texture has little effect on the model’s predictions. In contrast, applying SAR significantly increases the affected region on the depth map (high $\mathcal{E}_{\text{cover}}$), while omitting the merging loss results in a clearly visible vehicle contour (low $\mathcal{E}_{\text{blend}}$).
@@
To assess the real-world impact of our attack, we conduct system-level experiments on the open-source autonomous driving platform Baidu Apollo. In each scenario, the adversarial vehicle is placed ahead of the victim AV to simulate realistic traffic. We select representative scenes and feed perception outputs, after decision-level fusion (zhu2024malicious), into Apollo’s planning module under both benign and adversarial conditions. In the benign case, the AV detects the vehicle ahead, slows down, and initiates a lane change to avoid collision. In contrast, under attack, the adversarial texture disrupts scene understanding, causing the AV to maintain its speed and lane, ultimately resulting in a rear-end collision. As shown in Figure 7, the AV safely changes lane at a 10 m distance in the benign setting. However, in the adversarial case, it fails to respond and collides with the target in both tests. These results highlight the severe safety risks posed by stereo-consistent physical attacks.
@@ Acknowledgments
This research was supported by the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-GC-2023-006).
Supplementary Material
Figure 9: Depth maps comparison between original and rendered images.
Figure 10: Comparison between existing PAEs and our 3D texture-enabled merging attack PAEs from side view.
Figure 11: Comparison between existing PAEs and our 3D texture-enabled merging attack PAEs from back view.
Figure 12: Attack against SM-BDE under different external environment variations. (a) Merging attack. (b) Appearing attack
Stereo-Aligned 3D Rendering
To show the effectiveness of our rendering module, we randomly selected 600 pairs of stereo images from KITTI2014 (geiger2012we) with target vehicles. We utilize SMOKE (liu2020smoke) to retrieve the 3D vehicle bounding boxes and categorized them by box size (Extra-small, Small, Medium, Large, Extra-Large) and yaw angle into eight discrete ranges ( 0, 45, 90, 135, 180, 225, 275, 315). Finally, Stereo-Aligned 3D Rendering is used to project the same 3D vehicle mesh on these images based on the outputs of detection model.
Figure 9 presents five example scenes with bounding boxes of different sizes and yaw angles. We can visually see that the rendered vehicles precisely align with the contour of the original vehicles. The depth maps generated by different SM models (chang2018; zhang2019ga; lipson2021raft; li2022practical) for the original and 3D stereo-aligned images are almost identical. Table 6 and Table 7 show that, regardless of changes in SM model, bounding box size and yaw angle, the depth shift between the two types of depth maps consistently remains within the range of $(0,3)$.
Size Extra-small Small Medium Large Extra-Large
Amount 138 130 120 114 98
PSMNet 0.42 0.55 0.72 1.31 2.90
GANet 0.52 0.70 1.03 1.62 2.83
RAFT 0.41 0.63 0.99 1.56 2.53
CRE 0.31 0.42 0.82 1.57 2.61
Table 6: Depth shift $\mathcal{E}_{\text{shift}}\downarrow$ between the original and rendered images (sizes).
Angles() 0 45 90 135 180 225 270 315
Amount 83 68 64 82 85 66 67 85
PSMNet 0.78 1.68 1.76 0.69 0.71 1.51 1.80 0.42
GANet 0.72 1.21 1.13 0.50 0.66 1.63 1.54 0.39
RAFT 0.61 1.52 1.52 0.49 0.57 1.67 1.62 0.40
CRE 0.57 1.72 1.83 0.61 0.55 1.81 1.93 0.51
Table 7: Depth shift $\mathcal{E}_{\text{shift}}\downarrow$ between the original and rendered images (yaw angles).
Additional Experiments Results
Figure 10 and Figure 11 shows additional result comparisons between existing PAEs and our 3D texture-enabled PAEs from different angles. (a) Benign texture; (b) Random texture; (c), (d) 2D adversarial patches from (liu2024physical) and (cheng2021towards), respectively, applied to 3D vehicle. (e) MDE Adversarial texture from (zheng2024physical); (f) Our 3D merging attack adversarial texture. We can see that our 3D-enable adversarial texture remains effective from different viewing angles, while other methods consistently fail to produce any noticeable impact on the depth outputs.
We conducted additional test attacks under different weather and lighting conditions with default experimental settings. From the results shown in Table 8, we can see that different environmental conditions have minor impact on the effectiveness of the attack. The depth map outputs are also shown in Figure 12.
Weathers Appear Merge
$\mathcal{E}_{\text{shift}}\uparrow$ $\mathcal{E}_{\text{cover}}\uparrow$ $\mathcal{E}_{\text{blend}}\downarrow$ $\mathcal{E}_{\text{cover}}\uparrow$
Morning 18.83 0.951 0.058 0.512
Midday 16.34 0.836 0.053 0.590
Sunset 14.75 0.617 0.063 0.558
Night 15.38 0.570 0.065 0.596
Foggy 16.69 0.732 0.060 0.588
Rainy 15.98 0.609 0.069 0.634
Table 8: Test results of our Appearing attack and Merging attack in different environment conditions.

### v4

@@ Introduction
Figure 1: Comparison between existing PAEs and our 3D texture-enabled merging attack PAEs in CARLA simulator (Dosovitskiy et al. 2017). (a) Benign right and left camera images with the corresponding depth map; (b) Random texture; (c), (d) 2D adversarial patches from (Liu et al. 2024) and (Cheng and others 2021), respectively, applied to 3D vehicle. (e) MDE Adversarial texture from (Zheng et al. 2024); (f) Our 3D merging attack adversarial texture.
@@
Recent studies have revealed the vulnerabilities of deep neural perception models for Autonomous Driving (AD) against physical adversarial attacks (Lovisotto et al. 2021; Zhu et al. 2021b; Cao et al. 2021). These attacks rely on physically deployable artifacts, also known as Physical Adversarial Examples (PAEs) (Eykholt et al. 2018; Thys and others 2019), to affect the vision features captured by the neural networks and then manipulate their predictions. Currently, most PAEs are in the form of 2D patches that affect only a limited fraction of an object, and they often target only Monocular Depth Estimation (MDE) models (Cheng et al. 2022; Guesmi et al. 2023; Wong et al. 2021; Cheng and others 2021). Consequently, these threats to AD might be readily bypassed in real-world 3D scenarios with the support of stereo-based Binocular Depth Estimation (BDE) (Xu and Zhang 2020; Wang et al. 2019; Li et al. 2019).
@@
In fact, attacks on BDE driven by Stereo Matching (SM), a depth estimation method leveraging visual disparity, remain very limited. Sharply different from the PAEs for MDE, attacking SM models presents unique challenges due to its reliance on the physical 3D disparity between two images, rather than only on the 2D image pixels captured by a single camera; it is this uniqueness that invalidates most of MDE attacks under SM-BDE. Although a few proposals (Liu et al. 2024; Cheng and others 2021) have explored attacks on stereo matching, they mainly rely on 2D patches overlaid in image space, which not only limits their effectiveness to narrow viewpoints, but also breaks the physical disparity consistency required by real-world systems, rendering them ineffective under physical deployment. Meanwhile, SM-BDE is increasingly adopted in AD, with dedicated datasets available to provide stereo data for training SM-BDE models (Geiger et al. 2012; Mayer et al. 2016). Therefore, it makes sense to consider full-scale 3D PAEs that are effective against SM-BDE, because only such PAEs can substantiate their realistic threat to real-world AD. To generate such PAE, one would need a global 3D camouflage texture to cover the whole surface of a target object, because such a texture can significantly increase the affected range within camera views and thus achieve a full-scale adversarial influence on the target.
@@
Adapting PAEs to SM-BDE may also enable more meaningful attacks. Existing attacks on depth estimation primarily focused on hiding objects by pushing their estimated depth to infinity (Guesmi et al. 2023; Zheng et al. 2024; Cheng et al. 2022). As a result, such attacks fail to really hide any objects, as the contour of the object remains, totally exposing the attackers’ intention. A more meaningful attack, as one may intuitively expect, should involves adjustable factors for depth fine-tuning, so that attackers can completely adapt the depth of different parts of the target object (including contours) into the surroundings, which we refer to as merging attack. Nonetheless, realizing the 3D PAE faces a great challenge under SM-BDE, as each pair of rendered images must precisely align with the physical geometry of stereo cameras in the real world. In addition to making PAEs viable under SM-BDE, a qualified design should also remain robust to varying viewpoints (e.g., position and heading) of the SM camera pair, as well as to diversified external environments (e.g., lighting and weather conditions).
@@
Additionally, our PAE optimization accounts for diverse environmental variations, including changes in distances, rotations lighting, and weather conditions. Figure 1 presents our merging attack PAEs compared to prior works (Liu et al. 2024; Cheng and others 2021; Zheng et al. 2024). The results demonstrate that, under real-world 3D stereo matching constraints, only our method remains effective. Our main contributions are summarized as follows:
@@
SM-BDE has received significant attention for its reliability and efficiency in depth estimation for AD (Baidu 2025; Waymo 2025; Mobileye 2025). Early works (Zagoruyko and Komodakis 2015; Zbontar and LeCun 2016) leverage deep networks for separate feature extraction and matching. Subsequent advances include PSMNet (Chang and Chen 2018), which incorporates pyramid pooling and 3D CNNs; GA-Net (Zhang et al. 2019), which introduces attention-based cost aggregation; and RAFT-Stereo (Lipson et al. 2021), which uses recurrent units for multi-level refinement. Recent works explore cascaded recurrent networks, such as CREStereo (Li et al. 2022), and focus on robust generalization and arbitrary-scale disparity, as demonstrated by AnyStereo (Liang and Li 2024). Given the rapid development of SM-BDE in AD, addressing its security has become critical.
@@
Adversarial attacks exploit vulnerabilities in DNN by introducing crafted perturbations to input data, causing the model to make incorrect predictions (Goodfellow and others 2014; Madry et al. 2017). PAEs have shown real-world effectiveness (Kurakin and others 2018; Eykholt et al. 2018), sparking broad interest in safety-critical domains such as AD. They have been studied extensively across key AD tasks, including 3D and 2D object detection (Tu et al. 2020; Abdelfattah et al. 2021; Cao et al. 2021), object tracking (Schmalfuss et al. 2022; Ranjan et al. 2019), object segmentation (Zhu et al. 2021a; Xie et al. 2017), and MDE (Cheng et al. 2022; Guesmi et al. 2023; Wong et al. 2021). Recent works (Liu et al. 2024; Cheng and others 2021) also explored physical attacks on SM-BDE using 2D PAE but oversimplify stereo image relationships by assuming a direct shift between left and right views, neglecting stereo viewpoint differences. While this assumption may be efficient for digital 2D patch attacks, it fails to generalize to the complexities of real-world 3D scenarios, as shown in Figure 1. However, our 3D camouflage texture as PAE can conform to the parametric constraints of stereo cameras and achieve state-of-the-art attack effectiveness across diverse viewpoints on SM models.
@@
To ensure that the adversarial texture is both physically printable and visually smooth, we further incorporate two regularization terms: a total variation (TV) loss ($\mathcal{L}_{\text{tv}}$(Mahendran and Vedaldi 2015) that suppresses high-frequency noise, and a non-printability score (NPS) loss ($\mathcal{L}_{\text{nps}}$(Sharif et al. 2016) that encourages the use of printer-reproducible colors. The final loss function is:
@@
We train our adversarial textures using the KITTI2015 dataset (Geiger et al. 2012), a widely adopted benchmark for AD that provides stereo image pairs. The training samples cover diverse vehicle poses, scales, and backgrounds, promoting generalization across real-world scenes. To enhance the robustness of adversarial textures under diverse environmental conditions, we adopt Expectation over Transformation (EoT) (Athalye et al. 2018) during training. Specifically, we randomly perturb the position of a point light source within the range [3,3][-3,3] meters along each axis, and uniformly sample ambient light intensity from $[0.3,0.9]$. To simulate challenging weather such as rain and fog, we additionally inject random Gaussian noise into the rendered images during training. For boundary depth extraction, the kernel size $p_{k}$ is set to $40$. We optimize the adversarial texture using a differentiable renderer and train for 100 epochs with the Adam optimizer (Kingma 2014), with an initial learning rate of 0.01, scheduled via cosine decay to a minimum of 1e41e-4. The loss coefficients are set to $\alpha=5$ and $\beta=0.1$. All experiments are performed on a single NVIDIA RTX A5000 GPU.
@@
We compare our method with the two most relevant physical attacks designed against SM-BDE: PASM (Liu et al. 2024) and Adv-DM (Cheng and others 2021). In addition, we include two baselines textures: benign (neutral textures) and random (random noise). We evaluate all textures on five representative SM-BDE models: PSMNet (Chang and Chen 2018), GA-Net (Zhang et al. 2019), RAFT-Stereo (Lipson et al. 2021), CREStereo (Li et al. 2022), and AnyStereo (Liang and Li 2024), covering diverse cost volume constructions and model designs.
@@
We evaluate the trained textures in the CARLA simulator (Dosovitskiy et al. 2017) using a world-aligned stereo setup. Adversarial vehicles are placed at distances ranging from 3 to 30 meters and orientations between $0^{\circ}$ and $360^{\circ}$, under varying lighting and weather conditions in the “Town 10” map. Table 1 summarizes the evaluation results of our merging attack in comparison with state-of-the-art physical attacks across five representative SM-BDE models. All attacks are tested under identical camera baselines and environmental conditions in the CARLA simulator. Our adversarial texture consistently achieves a higher $\mathcal{E}_{\text{cover}}$, indicating a broader impact on predicted depth. This can be attributed to the larger effective surface coverage and better generalization across object orientations. Moreover, our approach attains the lowest $\mathcal{E}_{\text{blend}}$ among all methods, suggesting superior blending into the surrounding environment and reduced detectability in depth predictions, both key objectives of hiding attacks. Importantly, when integrated into a full commercial AD pipeline – Apollo, our attack leads to the highest $\mathcal{E}_{\text{sys}}$, indicating a significant increase in missed detection and path planning failures. This highlights the safety risk posed by our physically realizable texture.
@@
We conduct an ablation study to evaluate the effectiveness of the Stereo-Aligned 3D Rendering (SAR) module and the merging attack mechanism. We compare adversarial textures generated under four configurations: no module (None), SAR only, merging only, and the full method with both components. In the merging only setting, we define the main loss $\mathcal{L}_{\text{main}}$ as $\mathcal{L}_{\text{merge}}$ and reuse the same rendering parameters $k$ from the left view for the right. For None and SAR only, we replace $\mathcal{L}_{\text{merge}}$ with a standard hiding loss (Zheng et al. 2024). As shown in Table 4, removing the physical constraint $\mathcal{L}_{\text{nps}}$ slightly improves performance due to the relaxation of printability limits. However, without enforcing stereo consistency, the texture has little effect on the model’s predictions. In contrast, applying SAR significantly increases the affected region on the depth map (high $\mathcal{E}_{\text{cover}}$), while omitting the merging loss results in a clearly visible vehicle contour (low $\mathcal{E}_{\text{blend}}$).
@@
To assess the real-world impact of our attack, we conduct system-level experiments on the open-source autonomous driving platform Baidu Apollo. In each scenario, the adversarial vehicle is placed ahead of the victim AV to simulate realistic traffic. We select representative scenes and feed perception outputs, after decision-level fusion (Zhu et al. 2024), into Apollo’s planning module under both benign and adversarial conditions. In the benign case, the AV detects the vehicle ahead, slows down, and initiates a lane change to avoid collision. In contrast, under attack, the adversarial texture disrupts scene understanding, causing the AV to maintain its speed and lane, ultimately resulting in a rear-end collision. As shown in Figure 7, the AV safely changes lane at a 10 m distance in the benign setting. However, in the adversarial case, it fails to respond and collides with the target in both tests. These results highlight the severe safety risks posed by stereo-consistent physical attacks.
@@ Acknowledgments
This research/project is supported by the National Research Foundation, Singapore, under its AI Singapore Programme (AISG Award No: AISG4-GC-2023-006-1B).